# A Turing Test for Artificial Nets devoted to model Human Vision

### v1

@@ 1.3 Our Proposal: An Easy-To-Use Turing-Test from Vision Science
In our proposal, we suggest two approaches for testing DNNs: one linear and one nonlinear, as shown in the table [^3] The graphical quality of tables, figures and diagrams in this preprint will be updated in the journal version. in Fig.1. For color processing, we first test DNNs from a linear perspective, focusing on spectral sensitivities, and then from a nonlinear perspective to examine adaptation, nonlinear responses, and color illusions. Similarly, for texture processing, we begin by testing the linear aspects, such as receptive fields and contrast sensitivities, and then explore nonlinear responses, including masking effects and contrast illusions. This dual approach provides a comprehensive framework for evaluating how DNNs replicate the complexity of visual processing observed in the human brain.
@@ 4 Conclusions
Authors thank Prof. Raul Santos for his invitation to participate in the Workshop on Evaluating Artificial Intelligence (Univ. of Bristol 2022). This work was partially funded by MICIIN projects obtained 2020 and 2021 by J. Malo and V. Laparra.

### v2

@@ 1.3 Our Proposal: An Easy-To-Use Turing-Test from Vision Science
In our proposal, we suggest two approaches for testing DNNs: one linear and one nonlinear, as shown in the table [^3] The graphical quality of tables, figures and diagrams in this preprint will be updated in the journal version. in Fig.1. For color processing, we first test DNNs from a linear perspective, focusing on spectral sensitivities, and then from a nonlinear perspective to examine adaptation, nonlinear responses, and color illusions. Similarly, for texture processing, we begin by testing the linear aspects, such as receptive fields and contrast sensitivities, and then explore nonlinear responses, including masking effects and contrast illusions [^4] The 2024 code for stimuli generation is here: http://isp.uv.es/docs/TheDecalogueCode2024.zip. This dual approach provides a comprehensive framework for evaluating how DNNs replicate the complexity of visual processing observed in the human brain.
@@ 4 Conclusions
Authors thank Prof. Raul Santos for his invitation to participate in the Workshop on Evaluating Artificial Intelligence (Univ. of Bristol 2022). This work was partially funded by MICIIN projects obtained 2020 and 2021 by J. Malo and V. Laparra. The latest version of the code to generate the stimuli (by 2024) was possible due to the grant of the BBVA Foundations of Science program: Mathematics, Statistics, Computational Sciences and Artificial Intelligence (VIS4NN).